OpenStudy (turingtest):

Probability: geometric random variables Let \(X\) be a geometric random variable with parameter \(p\). Find the probability that \(X\ge10\). Express your answer in terms of \(p\) so I said:\[P(X\ge10)=\sum_{k=10}^\infty p^{10}(1-p)^{k-10}=p^{10}\sum_{k=0}^\infty (1-p)^k=p^9\]but this is wrong. Where is my mistake?

OpenStudy (ybarrap):

$$ P(X\ge10)=\sum_{k=10}^\infty q^{k-1}p=p\sum_{k=9}^\infty q^{k}\\ =1-p\sum_{k=0}^8 q^{k}\\ =1-\cfrac{1-q^{9}}{1-q}\\ =1-\cfrac{1-(1-p)^9}{p}\\ =\cfrac{p-1-(1-p)^9}{p}\\ =\cfrac{-(1-p)^9}{p}-\cfrac{1}{p}+1 $$ Does this make sense?

OpenStudy (turingtest):

why yes, indeed it does. Thank you!

OpenStudy (ybarrap):

Oops, I found an error: $$ P(X\ge10)=\sum_{k=10}^\infty q^{k-1}p=p\sum_{k=9}^\infty q^{k}\\ =1-p\sum_{k=0}^8 q^{k}\\ =1-p\cfrac{1-q^{9}}{1-q}\\ =1-p\cfrac{1-q^{9}}{p}\\ =1-(1-q^9)\\ =q^9\\ =(1-p)^9 $$